OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (uri):

Tony and beinda age in total :56 belinda is =2t+8

OpenStudy (campbell_st):

let tony's age be t then belinda is 2t + 8 so you have a combined age of t + 2t + 8 = 56 solve for t

OpenStudy (campbell_st):

the solution to the above will give tony's age. Just subtract that from 56 to get belinda's age

OpenStudy (uri):

T (tony's age)+B(belinda's age) =56 B=56-T

OpenStudy (uri):

And 56-T=2t+8 So,56-8=2t+t 54=3t

OpenStudy (uri):

48=3t*

OpenStudy (uri):

So tony =16.

OpenStudy (uri):

Tony's age is 16 Belinda =? 56-16=40 So tony is 16 and belinda is 40.

OpenStudy (uri):

@Viceroy o_O?

OpenStudy (anonymous):

Thanks mates.

OpenStudy (anonymous):

can this be this way : let t be tony's and b be belinda's....... so t + b =56.......(i) b= 8 +2t.........(ii) solve simultenously the two equations make the 3rd eqn... 56-b=t..............(iii) substute the 3rd eqn into the 2nd eqn b=8 +2(56-b) b=8 +112 -2b 3b=120 b=40 so just work for the other value.. substute the known value of b into any of the eqns from..b=8+2t but b =40 40=8+2t 32=2t t=16 therefore...Tony's age is 16 and Belinda's age is 40.